OpenStudy (anonymous):

dimension of a rectangle are such that its length is 11in. more than width. if the length were doubled and if the width were decreased by 5 in. the area would be increased by 130 in.^2 what r the length and width of the rectangle

OpenStudy (anonymous):

put w = width, then \[l=w+11\] so initial area is \[w(w+11)\] if you double the length you get \[2l=2(w+11)\] degrease the width by 5 you get \[w-5\] and the area of the changed figure is \[2(w+11)(w-5)\] the second area is 130 larger than the first, so we can solve \[w(w+11)+130=2(w+11)(w-5)\]

OpenStudy (anonymous):

hard part was getting an equation to solve, it should be good from there. fairly sure this is correct because i think the problem was cooked to give a nice integer solution

OpenStudy (anonymous):

thank you